OpenStudy (anonymous):

Choose the correct simplification of the expression −6x^2(4x − 7x^2 − 2). 42x^4 + 24x^3 − 12x^2 −42x^4 + 24x^3 − 12x^2 −13x^4 − 2x^3 − 8x^2 42x^4 − 24x^3 + 12x^2

OpenStudy (anonymous):

got a guess for this one?

OpenStudy (anonymous):

do you know how to distribute?

OpenStudy (anonymous):

well the teacher told me its 3

OpenStudy (anonymous):

your teacher is wrong

OpenStudy (anonymous):

o.O thats somethin i dont hear everyday lol

OpenStudy (anonymous):

"do you know how to distribute?"

OpenStudy (anonymous):

uhmmm kinda. but ill probably get it wrong.

OpenStudy (anonymous):

Just the distribute −6x^2 and -2 that'll make it so you can narrow the answer down to the correct one, or you can do the rest too lol

OpenStudy (anonymous):

\[−6x^2(4x − 7x^2 − 2)\] \[=-6x^2\times 4x-6x^2\times (-7x^2)-6x^2\times (-2)\] is the first step

OpenStudy (anonymous):

as @gaara438125 said, the last term will be \(+12x^2\) which narrows the answer to one choice

OpenStudy (anonymous):

yup. thanks guys. but can u explain it fully ? i dont jus want the answers.

OpenStudy (anonymous):

and i have no idea who to give the fan/medal too lol

OpenStudy (anonymous):

satellite73 explained the process \[−6x^2(4x − 7x^2 − 2)\] \[(-6x^2\times 4x) + (-6x^2\times -7x^2) + (-6x^2\times -2)\]

OpenStudy (anonymous):

is there another way to do it ?